OpenStudy (anonymous):

(2/x-5)-(4/x+3)=(8/x^2-9)

OpenStudy (anonymous):

[2(x+3) -4(x-5)] / (x-5)(x+3) = 8/(x^2-9)

OpenStudy (anonymous):

(26-2x)/ (x-5)(x+3) = 8/ [(x-3)(x+3) ]

OpenStudy (anonymous):

multiply both sides by (x+3)

OpenStudy (anonymous):

(26-2x)/ (x-5) = 8/(x-3)

OpenStudy (anonymous):

Cross Multiply

OpenStudy (anonymous):

(26-2x)(x-3) = 8(x-5)

OpenStudy (anonymous):

-2x^2 -72x - 78 =8x -40

OpenStudy (anonymous):

2x^2+80x+38 =0 x^2 +40x +19 =0

OpenStudy (anonymous):

You have to solve the last thing by quadratic formula.

jhonyy9 (jhonyy9):

hello there is +32x not -72x

jhonyy9 (jhonyy9):

elecengineer please correct your calcules !!!

jhonyy9 (jhonyy9):

brittany check it there is not right

OpenStudy (anonymous):

yes jhonyy9, I also end up with equation x^2-12x+19=0

jhonyy9 (jhonyy9):

so than x_1,_2=(12+/-sqrt(144-76))/2 =(12+/-sqrt(68))/2 =6+/-sqrt(17)

OpenStudy (anonymous):

\[(\frac{2}{x}-5)-(\frac{4}{x}+3)=(\frac{8}{x^{2}}-9)\] \[\frac{2}{x}-5-\frac{4}{x}-3=\frac{8}{x^{2}}-9\] \[\frac{2}{x}-\frac{4}{x}-\frac{8}{x^2}=5+3-9\] \[-\frac{2}{x}-\frac{8}{x^{2}}=-1\] \[-2x-8=-x^{2}\] \[x^{2}-2x-8=0\]

OpenStudy (anonymous):

yes jhonyy9, I have the same http://www.wolframalpha.com/input/?i=2%2F%28x-5%29-4%2F%28x%2B3%29%3D8%2F%28x%5E2-9%29

jhonyy9 (jhonyy9):

ok but look it please iant have wrote one newly possibillity

jhonyy9 (jhonyy9):

how is it true ?

jhonyy9 (jhonyy9):

this will be need dessiding brittany

OpenStudy (anonymous):

PArenthesis were put in totaly different way, but I think 1st version is correct

OpenStudy (anonymous):

lol britany which is it?

jhonyy9 (jhonyy9):

ok IanT but you need continue again till the end - till solutions

OpenStudy (anonymous):

It's late at night. She is sleeoing already :-)). Good night to everybody!

jhonyy9 (jhonyy9):

where is this ? because here is morning 10,23

OpenStudy (anonymous):

\[(x+2)(x-4)=0\] \[x+2=0\] or \[x-4=0\] \[x=-2,4\]

jhonyy9 (jhonyy9):

yes x_1= -2 and x_2=4

OpenStudy (anonymous):

It was late at night in US.